A 77.0 mL aliquot of a 1.70 M solution is diluted to a total volume of 228 mL. A 114 mL portion of that solution is diluted by adding 125 mL of water. What is the final concentration? Assume the volumes are additive.
1
Expert's answer
2019-10-21T07:16:37-0400

n=Cm x V = 0.077 x 1.7 = 0.1309 moles.

0.1309/2= 0.0645 .

Cm=0.0645/0.239 = 0.2698 M.